OpenStudy (anonymous):

Which expression is the first step you take when finding the quotient of five over seven and three over two? five over seven times three over two five over seven times two over three seven over five times two over three seven over five times three over two

OpenStudy (anonymous):

Get new questions pls D:

OpenStudy (blackops2luvr):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge\rlap{\qquad\qquad\mathbf{\;\,Gold~Stars}}\bbox[5pt,gold,border:2px solid red]{\qquad\qquad\color{red}{\mathbf{Gold~Stars}}\qquad\qquad\quad}\\ \Huge\rlap{\mathbf{\,\;For~Everyone!}}\bbox[5pt,gold,border:2px solid red]{\color{red}{\mathbf{For~Everyone!}}}\\ \Huge \begin{matrix} \color{gold}{★★★}\\ \color{gold}{★★}\\ \color{gold}{★★★} \end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (anonymous):

changeh ueh pic >:|

OpenStudy (blackops2luvr):

y

OpenStudy (anonymous):

Objection

OpenStudy (anonymous):

Zambie whats your guess ?

OpenStudy (anonymous):

uhhh, ummm

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

"Finding the quotient" simply means dividing the first by the second. So, 5/7 divided by 3/2 To divide fractions, multiply the first by the inverse of the second. So, 5/7 times 2/3 (five over seven times 2 over three) I hope this helps you understand the problem!!!!! _NOVA (GODGAVEMEYOU)

OpenStudy (anonymous):

0-o no, it doesn't

OpenStudy (anonymous):

bo2 luver -_-

OpenStudy (anonymous):

Finding the quotient is dividing but if you are dividing two fractions then the second fraction is always flipped over. So the answer to you question is 5/7 * 2/3

OpenStudy (anonymous):

thats my pro, pic

OpenStudy (anonymous):

um, ok,

OpenStudy (anonymous):

not good at dividing fractions :|

OpenStudy (anonymous):

You'll understand it easier one fay tag me in any of your questions ill try to help but I think we're in the same grade

OpenStudy (anonymous):

wut grade?

OpenStudy (anonymous):

8thgrade

OpenStudy (anonymous):

no, 7th

OpenStudy (blackops2luvr):

ello b

OpenStudy (anonymous):

Oh well I am in a little easier math but ok just tag me!

OpenStudy (blackops2luvr):

zambie is my pic look like something to chu?